OpenStudy (anonymous):

If Quinn tosses a fair coin and then rolls a fair number cube labeled 1 through 6, what is the probability of tossing heads followed by rolling a number less than 5?

OpenStudy (anonymous):

A 3 B 2/3 C 1/3

OpenStudy (anonymous):

fist determine the probability of having a head, and probability of having number that is less than 5. 2nd identify are these two events independent from each other or not. 3rd if they are independent from each other P(A and B) = P(A)P(B)

Directrix (directrix):

The probability of tossing a head on one toss of a fair coin is 1/2. One represents the number of desirable outcomes and 2 the number of possible outcomes. The possibilities when tossing a fair number cube are 1, 2, 3, 4, 5, and 6. To roll a number smaller than 5 would mean to toss a 1, 2, 3, or 4. The probability of tossing (1, 2, 3, or 4) = 4/6. The event for which you want to know the probability is tossing a head AND tossing a digit smaller than 5). ---------- Events A and B are said to be Independent. That is, the outcome of the first toss has no effect on the outcome of the second toss. The probability of the simultaneous occurrences of two independent events is the product of the probabilities of each event: P(A and B) = P(A) . P(B) ---------------- Multiply the two probabilities of your problem to get the probability of the total event. Read more about this at the following link: http://library.thinkquest.org/11506/prules.html